# Chest X-ray Foundation Model with Global and Local Representations Integration

### v1

@@ 5.2 Opportunistic predictive power
Beyond thoracic disease detection tasks, we evaluated ChesXFound’s generalizability in opportunistic CXR interpretation. For this purpose, we requested access to the CXR arm of the PLCO trial [46], which includes digitally scanned CXR films and up to 25-year morality follow-up data. Using this dataset, we investigated CheXFound’s predictive capability for cardiovascular disease (CVD) risk and all-cause mortality estimation. We used CheXFound with ViT-L pretrained on CXR-1M in this experiment and compared CheXFound against two vision foundation models (RAD-DINO, EVA-X) and the end-to-end trained model, ConvNeXt.
@@
CXR-1M ViT-Base 0.219(.215–.223) 0.815(.811–.819) 0.632(.583–.679) 0.877(.853–.896) 0.957(.924–.979) 0.947(.925–.963) 0.923(.788–1.0) 0.927(.818–1.0) 0.908(.806–.955) 0.845(.696–.943)
ViT-Large 0.252(.247–.258) 0.830(.826–.834) 0.679(.630–.727) 0.908(.894–.921) 0.983(.960–.996) 0.978(.951–.995) 1.000(1.0–1.0) 1.000(1.0–1.0) 0.986(.956–1.0) 0.975(.931–1.0)
@@ 5.3 Scalability of self-supervised vision encoders
The scaling capabilities of the self-supervised vision encoders depends on both the model size and the pretraining data size and diversity [21, 26]. To analyze the scaling trends, we pretrained CheXFound across a range of data scales, with CXR-1M and its two subsets CXR-744K and CXR-207K. We also evaluated the impact of model scale by using ViT-Base (ViT-B) and ViT-Large (ViT-L) as the backbones.
Our results in Table 5 demonstrate that CheXFound benefits form both data and model scaling. Increasing the pretraining data from CXR-207K to CXR-1M with a ViT-L backbone leads to significant AUROC improvements of 3.5% ($p<0.001$) on CXR-LT 24, 3.4% ($p<0.001$) on CheXpert, 3.8% ($p<0.001$) on Shenzhen, 6.8% ($p<0.001$) on Montgomery, and 30.7% ($p<0.001$) on JSRT. We observed similar trends when using ViT-B, with performance also improving as we scale up the pretraining data from CXR-207K to CXR-1M. In addition, CheXFound with ViT-L consistently outperformed the ViT-B architecture across different data sizes. These results align with previous studies on scaling ViT models [21, 12, 26].
@@ 6.1 Interpretation of disease-specific local features
The interpretabliliy of an artificial intelligence model is crucial to its medical applications. In this study, we trained GLoRI on top of the frozen foundation model. GLoRI inherently provides interpretable attention maps for each pathology. We visualized the attention maps for a selection of five pathologies of 25 CXRs and aligned these attention maps to an anchor CXR via affine registration [^1] We apply affine registration using the SimpleElastix library: https://simpleelastix.readthedocs.io. to provide a global perspective in Fig. 7. The attention maps contain precise localization of abnormalities in CXRs and the global attention maps cover the regions where the pathologies constantly occur. To be specific, edema refers to the accumulation of excess fluid and its abnormal regions often diffuse across lungs. This pattern is well captured by our attention maps as we observed in Fig. 7d that the critical regions in the individual attention maps scatter over the lungs and the global attention map covers extensive regions of both lungs. However, these attention maps have the limitations of covering partial regions of the abnormalities. For example, the attention maps for cardiomegaly only cover the heart on the left and right regions of the spine, and some maps for pleural effusion only cover the inferior boundaries of the lung while ignore the remaining abnormal regions.
@@ 6.2 Generalizability of foundation models
An important characteristic of CheXFound and other foundation models is their generalization capabilities to in-distribution and out-of-distribution downstream tasks. Compared with other encoders, we found that CheXFound achieved better performances on both in-distribution CXR-LT 24 and CheXpert datasets and out-of-distribution Shenzhen, Montgomery, and JSRT datasets. On the opportunistic CXR interpretation tasks on PLCO, CheXFound also achieved consistent and significant increases over comparison methods. CheXFound’s generalizability is attributed to the strong representation quality of frozen features learned via pretraining with large-scale, diverse CXRs. We also demonstrated CheXFound’s generalization capabilities on infrequent and underrepresented pathologies. CheXFound achieved significant increases over its comparions in classifying low-prevalence pathologies with lower than 1% occurrence frequencies, demonstrating its superior label efficiency. Although CheXFound with ViT-L achieved robust generalizability, our study did not evaluate the best-performing ViT-giant (ViT-g) architecture in DINOv2, a larger model with 1.1B parameters, which we expect to achieve better generalization performances in CXR interpretation, but it demands more pretraining data and computational resources. Overall, we demonstrated CheXFound’s robust generalization capabilities, which we believe can enable diverse downstream adaptations with improved label efficiency.
@@ 6.3 Conclusion
In summary, this work introduces CheXFound, a vision-centric foundation model pretrained via self-distillation on over one million unique CXRs. For downstream adaptations, we trained a GLoRI module on top of the frozen CheXFound, which combines disease-specific local features and global image features to improve the multilabel classification performance. CheXFound outperformed previous methods for classifying 40 disease findings on CXR-LT 24, demonstrating superior label efficiency on datasets with limited training labels, and strong generalization capabilities for opportunistic CXR interpretation on PLCO. The disease-specific local features extracted from CheXFound also carry strong interpretability as visuliazed by the associated attention maps. In our future work, we will continue to explore novel pretraining schemes to further improve the understanding of CXRs by these foundation models.

### v2

@@ 5.3 Opportunistic predictive power
Beyond thoracic disease detection tasks, we evaluated CheXFound’s generalizability in opportunistic CXR interpretation. For this purpose, we requested access to the CXR arm of the PLCO trial [57], which includes digitally scanned CXR films and up to 25-year mortality follow-up data. Using this dataset, we investigated CheXFound’s predictive capability for cardiovascular disease (CVD) risk and all-cause mortality estimation. We used CheXFound with ViT-L pretrained on CXR-987K in this experiment and compared CheXFound against two vision foundation models (RAD-DINO, EVA-X) and the end-to-end trained model, ConvNeXt.
@@
5.4 CXR foundation models for semantic segmentation
In addition to the classification tasks, we applied the pretrained CheXFound for semantic segmentation on CXRs. We included the VinDr-RibCXR [25] dataset which contains segmentation masks for 10 left ribs and 10 right ribs and the Montgomery dataset [53] which contains segmentation masks for left and right lungs. We trained a UPerNet decoder [58] on top of the frozen foundation models to perform segmentation. We compared CheXFound with other vision foundation models (EVA-X [11] and RAD-DINO [12]) and an end-to-end trained model (ConvNeXt [55]). As shown in Table 8, CheXFound achieved a Dice similarity coefficient of 0.793 on VinDr-RibCXR, outperforming the next best-performing model (ConvNeXt) by 7.1% ($p<0.001$). On Montgomery, CheXFound achieved a Dice similarity coefficient of 0.980, slightly surpassing the next best-performing model (RAD-DINO) by 0.2%. Such marginal performance gain over RAD-DINO on Montgomery likely reflects the relative simplicity of lung segmentation on this dataset, where all comparison methods achieved satisfactory performance. In contrast, the significant performance improvements in individual rib segmentation on VinDr-RibCXR demonstrates CheXFound’s superior capability to adapt to complicated anatomical structure segmentation tasks.
Table 8: Semantic segmentation results of the vision foundation models and end-to-end trained model. We trained a UPerNet [58] decoder on top of the frozen foundation model. We computed the mean and standard deviation of Dice similarity coefficients to evaluate segmentation performance. Values in bold indicate the best-performing results.
Models Frozen Backbone VinDr-RibCXR Montgomery
ConvNeXt [55] 0.710 $\pm$ 0.116 0.954 $\pm$ 0.047
EVA-X [11] 0.471 $\pm$ 0.108 0.956 $\pm$ 0.036
RAD-DINO [12] 0.684 $\pm$ 0.094 0.978 $\pm$ 0.006
CheXFound 0.793 $\pm$ 0.062 0.980 $\pm$ 0.005
Table 9: Benchmarking foundation models pretrained using different self-supervised methods on CXR-987K. We evaluated the downstream classification performance on CXR-LT 24. Values inside the parentheses are 95% confidence intervals. Values in bold indicate the best-performing results.
Methods Architectures AUPRC AUROC
Medical MAE [35] ViT-B/16 0.193(.189–.197) 0.772(.767–.779)
iBOT [9] ViT-L/16 0.212(.207–.217) 0.801(.796–.806)
EVA-X [11] ViT-B/16 0.168(.165–.171) 0.703(.698–.708)
RAD-DINO [12] ViT-B/14 0.185(.181–.189) 0.751(0.743–.759)
CheXFound ViT-L/16 0.265(.259–.271) 0.840(.836–.844)
@@ 5.4 CXR foundation models for semantic segmentation
CXR-987K ViT-Base 0.219(.215–.223) 0.815(.811–.819) 0.632(.583–.679) 0.877(.853–.896) 0.957(.924–.979) 0.947(.925–.963) 0.923(.788–1.0) 0.927(.818–1.0) 0.908(.806–.955) 0.845(.696–.943)
ViT-Large 0.265(.259–.271) 0.840(.836–.844) 0.679(.630–.727) 0.908(.894–.921) 0.983(.960–.996) 0.978(.951–.995) 1.000(1.0–1.0) 1.000(1.0–1.0) 0.986(.956–1.0) 0.975(.931–1.0)
5.5 Benchmarking self-supervised pretraining methods
We benckmarked various self-supervised pretraining methods on our curated CXR-987K dataset to evaluate their effectiveness in representation learning. By consistently using the CXR-987K dataset, we eliminated the influence of data heterogeneity in the pretraining stage. The self-supervised methods for comparison include Medical MAE [35], iBOT [9], EVA-X [11], and RAD-DINO [12]. Classification performance was evaluated on the CXR-LT 24 dataset. As shown in Table 9 below, CheXFound outperformed the next best-performing method (iBOT) by 3.9% ($p<0.001$) in AUROC. CheXFound achieved higher performance than iBOT owing to our advanced pretraining approach that utilized a low-resolution warm-up strategy and an optimized MIM loss weight on top of the original DINOv2. Medical MAE underperformed CheXFound because it lacks a discriminative objective which is critical for classification tasks. RAD-DINO, EVA-X, and Medical MAE were constrained by the shallower ViT-Base architecture, which encoded less effective representations than the ViT-Large architecture employed in our work.
@@ 5.6 Scalability of self-supervised vision encoders
The capabilities of the self-supervised vision encoders are affected by both the pretraining data scale and model scale [20, 28]. To analyze the scaling trends, we pretrained CheXFound across a range of data scales, including CXR-987K and its two subsets CXR-744K and CXR-207K. We also evaluated the impact of the model scale by using ViT-Base (ViT-B) and ViT-Large (ViT-L) as the backbones. We maintained consistent hyperparameters across all pretraining data sizes and model capacities. We trained CheXFound for 100 epochs (2,500 iterations per epoch) using the AdamW optimizer with a batch size of 14 and an initial learning rate of 2e-4. The resolution of input chest X-rays was set to 512$\times$512. We also applied a low-resolution warm-up strategy and a masked image modeling loss weight of 3, as determined by the ablation studies in Section 5.8.
Our results in Table 10 demonstrate that CheXFound benefits form both data and model scaling. Increasing the pretraining data from CXR-207K to CXR-987K with the ViT-L backbone leads to significant AUROC improvements of 3.5% ($p<0.001$) on CXR-LT 24, 3.4% ($p<0.001$) on CheXpert, 3.8% ($p<0.001$) on Shenzhen, 6.8% ($p<0.001$) on Montgomery, and 30.7% ($p<0.001$) on JSRT. We observed similar trends when using ViT-B, which showed the performance consistently improved as we increased the pretraining data scale from CXR-207K to CXR-987K. In addition, CheXFound with ViT-L consistently outperformed the ViT-B architecture across different pretraining data scales. These results align with previous studies on scaling ViT models [20, 12, 28].
@@ 5.7 Impact of CXR resolution
5.8 Ablation studies on pretraining
We evaluated the impact of various parameter initialization strategies on downstream performance. As shown in Fig. 7, random initialization and general domain initialization that utilized a checkpoint pretrained on natural images achieved AUROCs of 0.740 and 0.796, respectively, for classifying 40 diseases on CXR-LT 24. Incorporating a low-resolution warm-up strategy, which initializes pretraining at 512$\times$512 resolution with the checkpoint pretrained at 224$\times$224 resolution, increased the AUROC to 0.840, outperforming both the random initialization and general domain initialization strategies by large margins. We hypothesize that the degraded performance of direct high-resolution pretraining was due to the limited capability of the ViT architecture to capture extensive contextual information, which was mitigated by warming up the model with the low-resolution pretrained model that had already learned high-level contextual representations.
We further evaluated the impact of pretraining loss weights. The original DINOv2 settings employ balanced weights of 1:1 for the [CLS] token alignment and MIM losses for representation learning with natural images, where local patches often contain sufficient information to infer the global view. In contrast, medical images are characterized by subtle differences, with diagnostic cues often residing in nuanced local features. As such, we conducted multiple runs of pretraining with increased MIM loss weights to search for an optimal weight factor that adequately encourages local feature learning. As illustrated in Fig. 7, increasing MIM loss weights from 1 to 3 improved the AUROC from 0.822 to 0.840.
Overall, compared with the original DINOv2 settings that achieved an AUROC of 0.793 (Fig. 7) by training the model from scratch at 224$\times$224 resolution with a MIM loss weight of 1, our modifications to the initialization strategy for high-resolution pretraining and MIM loss weight delivered a performance gain of 4.7%.
Figure 7: Ablation studies on the impact of MIM loss weights and parameter initialization strategies on downstream classification. We evaluated the classification performance on the CXR-LT 24 dataset. Error bars indicate 95% confidence intervals.
Table 11: Ablation studies on the GLoRI architecture. We evaluated the classification performance on CXR-LT 24. We computed the mean AUPRC and AUROC values and their 95% confidence intervals over 1,000 bootstrapped samples. Values in bold indicate the best-performing results.
Architecture AUPRC $p$-value AUROC $p$-value
Linear probe 0.209(.204–.214) - 0.799(.794–.804) -
Attention pooler 0.246(.241–.252) - 0.826(.822–.830) -
+ Global representation 0.252(.247–.258) 0.023 0.830(.826–.834) 0.021
+ Adaptive Temperatures 0.262(.256–.268) $<0.001$ 0.835(.831–.839) $<0.001$
+ Pyramid Patch Merging 0.265(.259–.271) $<0.001$ 0.840(.836–.844) $<0.001$
Figure 8: Detailed results comparing GLoRI and linear probing for classifying different diseases. Diseases are sorted in descending order regarding their frequency on the dataset. We computed the mean and 95% confidence intervals of average precision using 1,000 bootstrapped samples. Error bars indicate 95% confidence intervals. Asterisks indicate statistical significance ($p<0.05$).”
5.9 Ablation studies on GLoRI head
We proposed GLoRI, a prediction head that incorporates both global and local image features for multilabel disease classification. Compared with prior attention pooling approaches [34, 50] that rely solely on an attention pooler with multi-head attention, GLoRI enhances classification performance by integrating an adaptive temperature module and a pyramid patch merging module into the multi-head attention mechanism and by incorporating the global image features. We evaluated GLoRI on the CXR-LT 24 dataset. As shown in Table 11, the attention pooler with a single multi-head attention layer achieved an AUROC of 0.826. This performance was improved by 0.4%, 0.5%, and 0.5% by incrementally incorporating the global representation, adaptive temperatures, and pyramid patch merging, respectively. Eventually, GLoRI achieved an AUROC of 0.84, which significantly improved the attention pooler by 1.4% ($p<0.001$, two-sided permutation test) and outperformed linear probe by 4.1% ($p<0.001$). We compared GLoRI with linear probe on detecting 40 diseases in Fig. 8. Overall, these improvements are attributed to GLoRI’s enriched representations, which integrate both the fine-grained and coarse-grained local features alongside the global image features.
@@ 6.1 Interpretation of disease-specific local features
The interpretability of an artificial intelligence model is crucial to its medical applications. In this study, we trained GLoRI on top of the frozen foundation model. GLoRI inherently provides interpretable attention maps for each pathology. We visualized the attention maps for a selection of five pathologies of 25 CXRs and aligned these attention maps to an anchor CXR via affine registration [^1] We apply affine registration using the SimpleElastix library: https://simpleelastix.readthedocs.io. to provide a global perspective in Fig. 9. The attention maps contain precise localization of abnormalities in CXRs and the global attention maps cover the regions where the pathologies constantly occur. To be specific, edema refers to the accumulation of excess fluid and its abnormal regions often diffuse across lungs. This pattern is well captured by our attention maps as we observed in Fig. 9 (d) that the critical regions in the individual attention maps scatter over the lungs and the global attention map covers extensive regions of both lungs. However, these attention maps have the limitations of covering partial regions of the abnormalities. For example, the attention maps for cardiomegaly only cover the heart on the left and right regions of the spine, and some maps for pleural effusion only cover the inferior boundaries of the lung while ignore the remaining abnormal regions.
@@ 6.2 Generalizability of foundation models
An important characteristic of CheXFound and other foundation models is their generalization capabilities to in-distribution and out-of-distribution downstream tasks. Compared with other encoders, we found that CheXFound achieved better performances on both in-distribution CXR-LT 24 and CheXpert datasets and out-of-distribution Shenzhen, Montgomery, and JSRT datasets. On the opportunistic CXR interpretation tasks on PLCO, CheXFound also achieved consistent and significant increases over comparison methods. CheXFound’s generalizability is attributed to the strong representation quality of frozen features learned via pretraining with large-scale, diverse CXRs. We also demonstrated CheXFound’s generalization capabilities on infrequent and underrepresented pathologies. CheXFound achieved significant increases over its comparisons in classifying low-prevalence pathologies with lower than 1% occurrence frequencies, demonstrating its superior label efficiency. Although CheXFound with ViT-L achieved robust generalizability, our study did not evaluate the best-performing ViT-giant (ViT-g) architecture in DINOv2, a larger model with 1.1B parameters, which we expect to achieve better generalization performances in CXR interpretation, but it demands more pretraining data and computational resources. Overall, we demonstrated CheXFound’s robust generalization capabilities, which we believe can enable diverse downstream adaptations with improved label efficiency.
@@ 6.3 Conclusion
In summary, this work introduces CheXFound, a vision-centric foundation model pretrained via self-distillation on over a large cohort of unique CXRs. For downstream adaptations, we trained a GLoRI head on top of the frozen CheXFound, which combined fine- and coarse-grained disease-specific local features and global image features to improve the multilabel classification performance. CheXFound outperformed previous methods for classifying 40 disease findings on CXR-LT 24. CheXFound demonstrated superior label efficiency on datasets with limited training labels and strong generalization capabilities to adapt to opportunistic CXR interpretation, malposition tube detection, and anatomical structure segmentation. The disease-specific local features extracted from CheXFound also manifested strong interpretability as characterized by the attention maps. In our future work, we will continue to explore novel pretraining schemes to further improve the understanding of CXRs by the foundation models.